Astronomy Report – October 23rd



Name: Jason Trump

Crew: 284

Date: Oct 23rd, 2023

Sky Conditions: Partly cloudy
Wind Conditions: Medium
Observation Start Time: 12:30pm
Observation End Time: 2:30pm
Summary: Crew visual observations which identified a large prominence near 7 o’clock. Solar imaging followed.
Objects Viewed: Sun
Images Submitted With This Report: Sun 231023 Prominence
Problems Encountered: Camera was left attached to telescope at start of mission. Confusion using Sharp Cap program because software update has caused user interface to differ from Quick Guide.

Journalist Report – October 23rd

Crew 284 has arrived at the MDRS as the 5th cohort of Spaceward Bound Utah: a science, sim, and education mission. We are a group of educators here to experience the Mars sim and explore avenues for bringing the science and sim experience back to our various learning environments: formal classrooms, science centers, education organizations, and more.

Our first day at the MDRS was not in sim, and instead focused on establishing protocols, crew rapport, and familiarization with the local region. Crew 284 completed the following activities:

– Completed facility training, including instruction on maintaining the HAB systems, science area, green HAB, and rover operations.
– Carried out a field exploration hike to learn more about the geology of the region around the MDRS.
– Utilized the Musk Observatory to observe current sun conditions.
– Visited the Hanksville-Burpee Dinosaur quarry to continue learning about the geologic conditions of the region.

Tomorrow we will enter Sim.

Jen Carver

Supplemental Operations Report – October 22nd

Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Ham radio has functional issues (it is transmitting a signal without any indications). It had been sent for repairs.

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 64 %; Avg 80.8%.
VDC Last 24 hours: Max 58.30V; Min 46.29V; Avg 51.09V.
Generator run time: 3846.5 hours.

Propane Readings:
Station Tank: 72%
Director Tank: 49%
Intern Tank: 69%
Generator Tank: 74%

Hab Static Tank – 530 gallons
GreenHab – 100 gallons
Outpost tank – 450 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: Yes.
Hours: 195.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: New battery cables for Spirit received. New parts are scheduled to be installed during crew 284 rotation.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Hanksville for supplies.
General notes and comments: Crew car new registration received, awaiting renewed insurance ID.

Summary of Internet: Nominal.

EVA suits and radios:
Suits: All nominal
Comms: All nominal

Campus wide inspection, if action taken, what and why: Nothing to report.
Summary of Hab Operations: Toilet tank cleaned and level sensors are functioning again. Remote temperature sensors for smart home systems are installed on the lower and upper deck. Ham radio has been taken for the repairs by Hope Lea and her father.
Summary of GreenHab Operations: Soil has been rehydrated. Interior reorganized by crew 283.
Summary of SciDome Operations: All nominal
Summary of Observatories Operations: All nominal.
Summary of RAM Operations: All nominal. Remote temperature sensors for smart home systems have been installed.
Summary of Outpost Operations: The Director’s trailer heater heater fixed (the control board replaced) and current work is nominal. A rat has been caught in the Intern trailer.
Summary of Health and Safety Issues: Nothing to report.

Copyright © The Mars Society. All rights reserved. | Main Site